OpenStudy (anonymous):

what's the inverse of y = 4x-2

OpenStudy (cwrw238):

find x in terms of y then replace the y by x - then you have the inverse

OpenStudy (anonymous):

so would it be y=1/4x-1/2 ?

OpenStudy (anonymous):

after adding 2, and dividing by 4, you are left with (y + 2)/4 = x if you replace x and y with each other now (x+2)/4 = y or simplified down to 1/2 + x/4 =y